Find the Hypotenuse of a right triangle whose legs are (a) 3, 4 (b) 5, 12 (c) 6, 8: (d) 7, 24 (e) 8, 15.
What do I do?

Answer by edjones(8007)

(a)
a^2+b^2=c^2
3^2+4^2
=9+16
=25
sqrt(25)=5 hypotenuse
.
Ed